OpenStudy (anonymous):

If susan subtracts 11 from one fourth of her number she gets 11.What is Susan's number?

OpenStudy (agl202):

x/4 -11 = 11 x/4 = 22 x=88

OpenStudy (ribhu):

let the no be x x/4 -11 =11 x/4 = 11+11 x/4 = 22 x = 88 this will be the number

OpenStudy (ribhu):

@Daina900 understood??